# Efficient Neural Combinatorial Optimization Solver for the Min-max Heterogeneous Capacitated Vehicle Routing Problem

### v2

@@ 4.3 Data Augment Method
SISR $\dagger$ 2.91 - 276 4.45 - 625
GA $\dagger$ 5.98 105.50 405 9.10 104.49 772
SA $\dagger$ 3.30 13.40 362 5.01 12.58 678
AM (ICLR’19) $\dagger$ (g.) 4.15 42.61 0.09 6.13 37.75 0.13
ET (AAAI’24) $\dagger$ (g.) 3.58 23.02 0.16 5.23 17.53 0.26
DPN (ICML’24) $\dagger$ (g.) 3.45 18.56 0.26 4.98 11.91 0.43
DRL (TCYB’22) $\dagger$ (g.) 3.60 23.71 0.25 5.38 20.90 0.43
2D-Ptr (AAMAS’24) (g.) 3.31 13.75 0.05 4.92 10.56 0.08
PARCO (NeurIPS’25) $\dagger$ (g.) 3.25 11.68 0.05 4.79 7.64 0.05
ECHO (ours) (g.) 3.05 4.71 0.06 4.61 3.61 0.09
AM (ICLR’19) $\dagger$ (s.) 3.63 24.74 0.14 5.58 25.39 0.28
ET (AAAI’24) $\dagger$ (s.) 3.33 14.43 0.22 4.98 11.91 0.40
DPN (ICML’24) $\dagger$ (s.) 3.20 9.97 0.38 4.79 7.64 0.78
DRL (TCYB’22) $\dagger$ (s.) 3.25 11.68 0.43 4.98 11.91 0.92
2D-Ptr (AAMAS’24) (s.) 3.09 6.19 0.09 4.68 5.17 0.18
PARCO (NeurIPS’25) $\dagger$ (s.) 3.06 5.15 0.07 4.58 2.92 0.09
ECHO (ours) (s.) 2.93 0.98 0.47 4.48 0.77 1.14
5 Experimental Results
5.1 Experimental Settings
Datasets and Hardware configuration: Following the prior studies [2dptr, drl], we assess the performance of ECHO on MMHCVRP instances with varying numbers of vehicles and nodes. Except for the cross-distribution generalization experiments in Section 5.3, the coordinates of customer nodes and depot node are uniformly sampled from the unit square $0,1]\times[0,1]$, and the demand of customer nodes are uniformly sampled from the set $\{1,2,\cdots,9\}$. For vehicles, capacities are uniformly sampled from $\{20,21,\cdots,40\}$, and speeds are uniformly sampled from the interval $[0.5,1]$. For each vehicle and node scale, the training dataset comprises the number of 1,280,000 instances, and the test dataset comprises 10,000 instances. All experiments are conducted on a computer equipped with an Intel(R) Core(TM) i9-12900K CPU and an NVIDIA RTX 4090 GPU (24GB).
Baselines and Evaluation Metrics: We compare ECHO against six NCO benchmarks, including AM [kool_attention_2019], DRL [drl], ET [et], DPN [dpn], 2D-Ptr [2dptr], and PARCO [parco]. Of these, DRL, 2D-Ptr, and PARCO are specifically designed for the MMHCVRP, ET and DPN target the Min-max TSP, and AM is tailored to TSP and CVRP. In addition, following the prior studies [2dptr, drl], we also include the heuristic algorithms GA [KARAKATIC2015519], SA [ILHAN2021100911], and the SOTA SISR [sisr] for comparison. The evaluation metrics include the objective value (Obj.), Gap, and computation time (Time), where Gap is defined as the relative difference in the ”objective value” between each method and the SISR heuristic. In addition, for NCO solvers, we report the results under two decoding strategies, namely Greedy and Sampling decoding strategies. With Greedy decoding, the solver always selects vehicle(s) and node(s) with the maximum probability at each time step. Whereas the sampling strategy randomly selects vehicle(s) and node(s) utilizing the learned probability, constructs the number of $k$ candidate solutions, and ultimately reports the best among them. In all experiments conducted in the paper, $k$ is set to 1280, following the precedent established by [kool_attention_2019].
@@ 5.3 Generalization Studies
To assess the cross-scale and cross-distribution generalization abilities of ECHO, we conduct extensive experiments in this subsection. Specifically, in Figs. 5 and 6, we present the generalization performance of ECHO versus 2D-Ptr across varying node and vehicle scales, respectively. The abbreviation M3-N60 denotes a configuration with three vehicles and 60 nodes. In Fig. 7, we assess the cross-distribution generalization performance of ECHO and 2D-Ptr. To generate clustered test instances, following the procedure of the prior study [invit], we sample three cluster centers for each instance, assign each node uniformly to one center, and perturb each node’s coordinates with Gaussian noise to induce the clustered distribution.
@@
Figure 7: Generalization performance of NCO solvers across different distribution patterns.
As shown in Fig. 5, ECHO achieves comparable cross-vehicle scale generalization performance to 2D-Ptr. Furthermore, even when trained exclusively on non‑target vehicle scales, ECHO still outperforms 2D‑Ptr trained directly on the target scale. For example, when trained on the M5‑N60 configuration, ECHO attains an objective value of 6.78 on M3‑N60, compared to 6.82 achieved by 2D‑Ptr trained and evaluated on M3‑N60. As shown in Fig. 6, ECHO also achieves high-level cross-node scale generalization ability compared to 2D-Ptr. As shown in Fig. 7, under the clustered distribution, ECHO consistently outperforms 2D‑Ptr across all vehicle and node scales. This advantage arises from our dual‑modality encoder, which more effectively captures local topological relationships among nodes, thereby reducing sensitivity to distribution shifts.
@@ 5.4 Ablation Studies
In this subsection, we conduct ablation studies to investigate the contributions of key ECHO design components. The performance is evaluated on MMHCVRP instances with the number of three vehicles and 60 nodes, and the results are presented in Table 6. Specifically, in w/o dual-modality encoder, the node encoder processes only node attributes, ignores edge attributes, and disables the cross‑attention operator used to fuse edge and node features. In w/o PFCA mechanism, the decoder omits historical vehicle information by disabling the PFCA mechanism. In w/o vehicle augment, only node data are augmented following the prior study [pomo], while vehicle data remain unchanged. As shown in Table 6, the experimental results demonstrate that each component contributes to the SOTA performance of ECHO.
@@
w/o PFCA mechanism 1.28
w/o vehicle augment 1.20
ECHO 1.01
6 Conclusion

### v3

@@ 5.1.2 Hyper-parameter
SISR 2.91 - 276 4.45 - 625
GA 5.98 105.50 405 9.10 104.49 772
SA 3.30 13.40 362 5.01 12.58 678
AM (ICLR’19) (g.) 4.15 42.61 0.09 6.13 37.75 0.13
ET (AAAI’24) (g.) 3.58 23.02 0.16 5.23 17.53 0.26
DPN (ICML’24) (g.) 3.45 18.56 0.26 4.98 11.91 0.43
DRL (TCYB’22) (g.) 3.60 23.71 0.25 5.38 20.90 0.43
2D-Ptr (AAMAS’24) (g.) 3.31 13.75 0.11 4.92 10.56 0.17
PARCO (NeurIPS’25) (g.) 3.25 11.68 0.05 4.79 7.64 0.05
ECHO (ours) (g.) 3.04 4.78 0.12 4.61 3.71 0.19
AM (ICLR’19) (s.) 3.63 24.74 0.14 5.58 25.39 0.28
ET (AAAI’24) (s.) 3.33 14.43 0.22 4.98 11.91 0.40
DPN (ICML’24) (s.) 3.20 9.97 0.38 4.79 7.64 0.78
DRL (TCYB’22) (s.) 3.25 11.68 0.43 4.98 11.91 0.92
2D-Ptr (AAMAS’24) (s.) 3.09 6.19 0.14 4.68 5.17 0.24
PARCO (NeurIPS’25) (s.) 3.06 5.15 0.07 4.58 2.92 0.09
ECHO (ours) (s.) 2.93 1.02 0.63 4.49 0.91 1.45
5.1.3 Baselines and Evaluation Metrics:
We compare ECHO against six NCO benchmarks, including AM (Kool et al., 2019), DRL (Li et al., 2022), ET (Son et al., 2024), DPN (Zheng et al., 2024), 2D-Ptr (Liu et al., 2024), and PARCO (Berto et al., 2025). Of these, DRL, 2D-Ptr, and PARCO are specifically designed for the MMHCVRP, ET and DPN target the Min-max TSP, and AM is tailored to TSP and CVRP. In addition, following the prior studies (Liu et al., 2024; Li et al., 2022), we also include the heuristic algorithms GA (Karakatič and Podgorelec, 2015), SA (İLHAN, 2021), and the SOTA SISR (Christiaens and Vanden Berghe, 2020) for comparison. The evaluation metrics include the average objective function values (Obj.), Gap, and inference time (Time) per problem instance, where Gap is defined as the relative difference in the ”objective value” between each method and the SISR heuristic. In addition, for NCO solvers, we report the results under two decoding strategies, namely Greedy and Sampling decoding strategies. With Greedy decoding, the solver always selects vehicle(s) and node(s) with the maximum probability at each time step. Whereas the sampling strategy randomly selects vehicle(s) and node(s) utilizing the learned probability, constructs the number of $k$ candidate solutions, and ultimately reports the best among them. In all experiments conducted in the paper, $k$ is set to 1280, following the precedent established by (Kool et al., 2019).
@@ 5.3 Generalization Studies
To assess the cross-scale and cross-distribution generalization abilities of ECHO, we conduct extensive experiments in this subsection. Specifically, in Figures. 5 and 6, we present the generalization performance of ECHO versus 2D-Ptr across varying node and vehicle scales, respectively. The abbreviation M3-N60 denotes a configuration with three vehicles and 60 nodes. Furthermore, in Table 8, we evaluate the generalization performance of ECHO and 2D-Ptr under three settings. Both ECHO and 2D-Ptr are trained on M7-N100 and then directly tested on these settings. Specifically, in Setting_1 and Setting_2, we evaluate the solvers on instances with nine vehicles and 120 customer nodes, as well as seven vehicles and 100 customer nodes, respectively, which lie outside the range of the training configurations. To introduce more extreme heterogeneity regimes, in Settting_3, the demands of customer nodes are sampled from a Gaussian distribution $\mathcal{N}(5,4)$ and then discretized to integers within the set $\{1,2,\cdots,9\}$, while vehicle capacities are uniformly sampled from $\{20,21,\cdots,80\}$. Finally, in Figures 7 and  8, we assess the generalization performance of ECHO and 2D-Ptr under Clustered and Explosion distribution patterns, respectively. To generate Clustered test instances, following the procedure of the prior study (Fang et al., 2024), we sample three cluster centers for each instance, assign each node uniformly to one center, and perturb each node’s coordinates with Gaussian noise to induce the clustered distribution. To generate Explosion test instances, following the procedure in the prior study (Fang et al., 2024), we first generate node coordinates from a uniform distribution. Then, a disc is defined by uniformly selecting a center on the board and a radius sampled uniformly from $[0.1,0.5]$. All nodes located inside this disc are subsequently displaced outward according to an exponential distribution with rate 10.
As shown in Figure 5, ECHO achieves comparable cross-vehicle scale generalization performance to 2D-Ptr. Furthermore, even when trained exclusively on non‑target vehicle scales, ECHO still outperforms 2D‑Ptr trained directly on the target scale. For example, when trained on the M5‑N60 configuration, ECHO attains an objective value of 6.78 on M3‑N60, compared to 6.82 achieved by 2D‑Ptr trained and evaluated on M3‑N60. As shown in Figure 6, ECHO also achieves high-level cross-node scale generalization ability compared to 2D-Ptr. As shown in Table 8, ECHO outperforms 2D-Ptr across all settings, even when evaluated on instances whose scales lie outside the training configurations and under more extreme heterogeneity regimes.
Table 8: Generalization performance of NCO solvers under Challenging Settings
Solver Obj. $\downarrow$
Setting_1 Setting_2 Setting_3
2D-Ptr (AAMAS’24) 4.27 6.80 3.54
ECHO (ours) 4.17 6.57 3.51
Figure 7: Generalization performance of NCO solvers under the Clustered distribution.
Figure 8: Generalization performance of NCO solvers under the Explosion distribution.
As shown in Figures 7 and 8, under the Clustered and Explosion distribution patterns, ECHO consistently outperforms 2D‑Ptr across all vehicle and node scales. This advantage arises from our dual‑modality encoder, which more effectively captures local topological relationships among nodes, thereby reducing sensitivity to distribution shifts.
@@ 5.4 Ablation Studies
In this subsection, we conduct ablation studies to investigate the contributions of key ECHO design components. The performance is evaluated on MMHCVRP instances with the number of three vehicles and 60 nodes, and the results are presented in Table 9. Specifically, in w/o dual-modality encoder, the node encoder processes only node attributes, ignores edge attributes, and disables the cross‑attention operator used to fuse edge and node features. In w/o PFCA mechanism, the decoder omits historical vehicle information by disabling the PFCA mechanism. In w/o vehicle augment, only node data are augmented following the prior study (Kwon et al., 2020), while vehicle data remain unchanged. As shown in Table 9, the experimental results demonstrate that each component contributes to the SOTA performance of ECHO.
@@
w/o PFCA mechanism 1.18
w/o vehicle augment 1.12
ECHO 1.00
5.5 Case Study
To illustrate the effectiveness of the proposed PFCA mechanism, we visualize the visiting trajectories produced by 2D-Ptr and ECHO on the same instance. As shown in Figure 9, ECHO tends to extend the route of the vehicle selected in the previous step, which can be attributed to the PFCA mechanism. In contrast, 2D-Ptr occasionally ignores the priority of the vehicle selected in the preceding step, leading to worse performance (Obj. = 5.04). For instance, after assigning the blue vehicle to visit the 31st customer at the 42nd step, 2D-Ptr does not select this vehicle again until the 57th step.
(a) 2D-Ptr (Obj. 5.04)
(b) ECHO (Obj. 4.79)
Figure 9: Comparison of decoding trajectories produced by 2D-Ptr and ECHO.
6 Discussions
From a macro perspective, ECHO and other NCO solvers designed for VRPs share a fundamental limitation in that their performance may degrade on real-world instances, because they are predominantly trained on synthetic data with predefined distributions that can be accurately modeled. In contrast, real-world VRP instances often exhibit highly uncertain and complex distribution patterns that are difficult to model accurately. Consequently, even models with strong generalization capabilities may experience inevitable performance degradation when deployed in real-world scenarios (Wu et al., 2024b). Moreover, ECHO assumes Euclidean distance when constructing routing costs, whereas real-world logistics systems typically involve more complex measures, e.g., Manhattan distance in urban grids (Son et al., 2026). Therefore, future research could focus on developing NCO solvers that incorporate real-world distance and cost metrics, thereby improving their adaptability to practical path planning scenarios.
7 Conclusions